OpenStudy (anonymous):

(please show how you did this! and maybe even explain?) Write the equation in Y intercept form: y-5=.75(or three fourths)(x-5)

OpenStudy (anonymous):

ok wait first replace x with zero

OpenStudy (anonymous):

o get y-5=-15/4

OpenStudy (anonymous):

now

OpenStudy (anonymous):

y=-15/4+5 y=5/4

OpenStudy (anonymous):

and y=-5/4 y intercept

OpenStudy (anonymous):

srry y=5/4

OpenStudy (anonymous):

u confused then ask

OpenStudy (anonymous):

welll theres answer choices and they are y=3/4x-5/4, y=-3/4x+5/4, y=3/4x+5/4, and y =3/4x-3/5

OpenStudy (anonymous):

m confused how can it be ....